OpenStudy (anonymous):

simplify (y^-7*y^-3)^-1

OpenStudy (anonymous):

\[(y^{-7} * 7^{-3})^{-1}\] ^^^^ that?

OpenStudy (anonymous):

that is the correct problem the way you set it up but I could not figure out how to set it up correctly

OpenStudy (anonymous):

(y^{-7} * 7^{-3})^{-1} in the equation bar

OpenStudy (anonymous):

ok so, do you know what we do with variables to powers when multiplied?

OpenStudy (anonymous):

subtract right

OpenStudy (anonymous):

that's divide, so if we multiply we add them

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

lol are you done?

OpenStudy (anonymous):

\[(1/(y^7*y^3))^(-1) \]\[\left(1\left/\left(y^{10}\right)\right.\right){}^{\wedge}(-1) \]\[\left(\frac{1}{y^{10}}\right){}^{\wedge}(-1) \]\[y^{10} \]